# 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for 3D Object Detection

### v1

@@ A.2 Future Work
Seg-Cluster (Wang et al., 2018) - - - - - - - - - -
Mask R-CNN (He et al., 2017) - - - - - - - - - -
SGPN (Wang et al., 2018) - - - - - - - - - -
3D-SIS (Hou et al., 2019) 5.1 42.2 50.1 31.8 15.1 1.4 0.0 1.4 0.0 0.0
3D-SIS (w/ RGB) (Hou et al., 2019) 5.7 50.3 52.6 55.4 22.0 10.9 0.0 13.2 0.0 0.0
VoteNet (Qi et al., 2019) 8.1 76.1 67.2 68.8 42.4 15.3 6.4 28.0 1.3 9.5
FCAF3D (Rukhovich et al., 2022a) 35.8 81.5 89.8 85.0 62.0 44.1 30.7 58.4 17.9 31.3
CAGroup3D (Wang et al., 2022a) 41.4 82.8 90.8 85.6 64.9 54.3 37.3 64.1 31.4 41.1
ImGeoNet (Tu et al., 2023) 15.8 74.8 46.5 45.7 39.9 8.0 2.9 32.9 0.3 7.9
CN-RMA (Shen et al., 2024a) 21.3 69.2 52.4 63.5 42.9 11.1 6.5 40.0 1.2 24.9
ImVoxelNet (Rukhovich et al., 2022b) 8.9 67.1 35.0 33.1 30.5 4.9 1.3 7.0 0.1 0.9
NeRF-Det (Xu et al., 2023) 12.0 68.4 47.8 58.3 42.8 7.1 3.0 31.3 1.6 11.6
@@
Seg-Cluster (Wang et al., 2018) - - - - - - - - -
Mask R-CNN (He et al., 2017) - - - - - - - - -
SGPN (Wang et al., 2018) - - - - - - - - -
3D-SIS (Hou et al., 2019) 13.7 0.0 2.7 3.0 56.8 8.7 28.5 2.6 14.6
3D-SIS (w/ RGB) (Hou et al., 2019) 23.6 2.6 24.5 0.8 71.8 8.9 56.4 6.9 22.5
VoteNet (Qi et al., 2019) 37.5 11.6 27.8 10.0 86.5 16.8 78.9 11.7 33.5
FCAF3D (Rukhovich et al., 2022a) 53.4 44.2 46.8 64.2 91.6 52.6 84.5 57.1 57.3
CAGroup3D (Wang et al., 2022a) 63.6 44.4 57.0 49.3 98.2 55.4 82.4 58.8 61.3
ImGeoNet (Tu et al., 2023) 43.9 4.3 24.0 2.0 68.8 24.5 61.7 17.4 28.9
CN-RMA (Shen et al., 2024a) 51.4 19.6 33.0 6.6 73.3 36.1 76.4 31.5 36.8
ImVoxelNet (Rukhovich et al., 2022b) 35.5 0.6 22.1 4.5 67.7 18.9 60.2 10.1 22.7
NeRF-Det (Xu et al., 2023) 46.0 5.8 26.0 1.6 69.0 25.5 55.8 21.1 29.7
@@
3DGS-DET (Our basic pipeline+BG+BS) 52.4 22.2 36.9 15.7 82.6 35.1 74.0 28.9 37.8 (+8.1)
@@
Figure 6: More qualitative comparison. Our methods identify more objects in the scene with better positional precision, highlighting the advantages of our approach over NeRF-Det (Xu et al., 2023). In this figure, the scene is represented using mesh to clearly display the boxes. Note that, Black and white boxes indicate predictions with incorrect categories, while boxes of other colors represent predictions with the correct category.
@@
Figure 7: More qualitative comparison. Our methods identify more objects in the scene with better positional precision, highlighting the advantages of our approach over NeRF-Det (Xu et al., 2023). In this figure, the scene is represented using mesh to clearly display the boxes. Note that, Black and white boxes indicate predictions with incorrect categories, while boxes of other colors represent predictions with the correct category.
@@
Figure 8: Analysis on the effect of Boundary Guidance. Gaussian blobs trained with Boundary Guidance exhibit clearer spatial distribution and more distinct differentiation between objects and background. Note that we visualize only the positions of the Gaussian blobs to highlight their spatial distribution, omitting other attributes.
@@
Figure 9: Rendered images from different views by 3DGS trained with Boundary Guidance. The category-specific boundaries are well rendered and exhibit multi-view stability, demonstrating that the 3D representation has successfully embedded the priors provided by Boundary Guidance.
@@
Figure 10: Rendered images from different views by 3DGS trained with Boundary Guidance. The category-specific boundaries are well rendered and exhibit multi-view stability, demonstrating that the 3D representation has successfully embedded the priors provided by Boundary Guidance.
@@
Figure 11: Analysis on the effect of Box-Focused Sampling. Box-Focused Sampling significantly retains more object blobs and reduces noisy background blobs. Note that we visualize only the positions of the Gaussian blobs to highlight their spatial distribution, omitting other attributes.

### v2

@@ 2.6 Efficiency analysis
VoteNet [33] 8.1 76.1 67.2 68.8 42.4 15.3 6.4 28.0 1.3 9.5
FCAF3D [37] 35.8 81.5 89.8 85.0 62.0 44.1 30.7 58.4 17.9 31.3
CAGroup3D [46] 41.4 82.8 90.8 85.6 64.9 54.3 37.3 64.1 31.4 41.1
ImGeoNet [45] 15.8 74.8 46.5 45.7 39.9 8.0 2.9 32.9 0.3 7.9
CN-RMA [41] 21.3 69.2 52.4 63.5 42.9 11.1 6.5 40.0 1.2 24.9
ImVoxelNet [38] 8.9 67.1 35.0 33.1 30.5 4.9 1.3 7.0 0.1 0.9
NeRF-Det [52] 12.0 68.4 47.8 58.3 42.8 7.1 3.0 31.3 1.6 11.6
NeRF-Det++ [16] - - - - - - - - -
@@
VoteNet [33] 37.5 11.6 27.8 10.0 86.5 16.8 78.9 11.7 33.5
FCAF3D [37] 53.4 44.2 46.8 64.2 91.6 52.6 84.5 57.1 57.3
CAGroup3D [46] 63.6 44.4 57.0 49.3 98.2 55.4 82.4 58.8 61.3
ImGeoNet [45] 43.9 4.3 24.0 2.0 68.8 24.5 61.7 17.4 28.9
CN-RMA [41] 51.4 19.6 33.0 6.6 73.3 36.1 76.4 31.5 36.8
ImVoxelNet [38] 35.5 0.6 22.1 4.5 67.7 18.9 60.2 10.1 22.7
NeRF-Det [52] 46.0 5.8 26.0 1.6 69.0 25.5 55.8 21.1 29.7
NeRF-Det++ [16] - - - - - - - - 30.0
@@
3DGS-DET (Our basic pipeline+BG+BS) 52.4 22.2 36.9 15.7 82.6 35.1 74.0 28.9 37.8 (+7.8)
Table 3: Comparison of mAP@0.5 on ScanNet. The first two blocks include methods that utilize non-view-synthesis representations: the first block consists of methods based on point clouds [33, 37, 46], while the second block focuses on multi-view image methods [45, 41, 38]. The third block includes methods that employ view-synthesis representations, including NeRF-based approaches and our 3DGS-based method. Among non-view-synthesis representations, our approach clearly outperforms methods that utilize multi-view images (the second block), demonstrating the superiority of our 3DGS reconstruction for detection with multi-view image inputs. In terms of view-synthesis representations, our 3DGS-DET significantly surpasses the NeRF-based method NeRF-Det++ [16] by 7.8 points. NeRF-Det++ [16] does not report per-category performance of mAP@0.5. We report its overall mAP@0.5 following the original paper [16].
3 Additional Main Results
3.1 Additional Quantitative Results
In Tab. 3, we compare the mAP@0.5 performance of various methods on the ScanNet dataset. The methods are grouped into three blocks based on their underlying representations. The first two blocks include methods that utilize non-view-synthesis representations: the first block consists of methods based on point clouds [33, 37, 46], while the second block focuses on multi-view image methods [45, 41, 38]. The third block includes methods that employ view-synthesis representations, including NeRF-based methods [52, 16] and our 3DGS-based method. Among non-view-synthesis representations, our method achieves clearly superior performance compared to methods utilizing multi-view images (the second block), demonstrating the effectiveness of our 3DGS reconstruction for detection tasks with multi-view image inputs. As for the view-synthesis representations, our 3DGS-DET outperforms the NeRF-based method NeRF-Det++ [16] by a significant margin of 7.8 points, further highlighting the advantages of our 3DGS-DET.
3.2 Additional Qualitative results
We provide more qualitative results in Fig. 1 and Fig. 2. As illustrated, our method detects more objects in the scene with higher positional accuracy than NeRF-Det [52], demonstrating the advantages of our method. Note that NeRF-Det++ [16] had not released pretrained models, code, or qualitative results on the official GitHub page by the time of our submission. Therefore, we could not provide a qualitative comparison with NeRF-Det++ here.
4 Failure Cases
We provide failure cases in Fig. 7. In scenes with severe occlusions, our method exhibits a small number of missed detections. These missed detections are caused by occlusions, which introduce inherent ambiguity in multi-view detection when key visual evidence is obscured.
5 Limitation and Future Work
While 3DGS-DET achieves significant improvements over strong alternative methods, several limitations remain for further exploration. As the first work to introduce 3DGS into indoor 3D Object Detection (3DOD), this work mainly focuses on the primary stage of this pipeline: empowering 3DGS for 3DOD. Extensive experiments demonstrate that our designs can lead to significant improvements. Beyond empowering the 3DGS representation, a subsequent detector specifically designed for 3DGS could hold promise in the future.
@@ 5 Limitation and Future Work
Figure 1: More qualitative comparisons. Our method detects more objects with better positional precision, highlighting the advantages of our approach over NeRF-Det [52]. In this figure, the scene is represented using mesh to clearly display the boxes. Note that black and white boxes indicate predictions with incorrect categories, while boxes of other colors represent predictions with the correct categories.
@@
Figure 2: More qualitative comparisons. Our method detects more objects with better positional precision, highlighting the advantages of our approach over NeRF-Det [52]. In this figure, the scene is represented using mesh to clearly display the boxes. Note that black and white boxes indicate predictions with incorrect categories, while boxes of other colors represent predictions with the correct categories.
@@
Figure 3: Analysis on the effect of Boundary Guidance. Gaussian blobs trained with Boundary Guidance exhibit clearer spatial distribution and more distinct differentiation between objects and background. Note that we visualize only the positions of the Gaussian blobs to highlight their spatial distribution, while omitting other attributes for clarity.
@@
Figure 4: Rendered images from different views by 3DGS trained with Boundary Guidance. The category-specific boundaries are clearly rendered and exhibit multi-view consistency, demonstrating that the 3D representation has successfully embedded the priors provided by Boundary Guidance.
@@
Figure 5: Rendered images from different views by 3DGS trained with Boundary Guidance. The category-specific boundaries are clearly rendered and exhibit multi-view consistency, demonstrating that the 3D representation has successfully embedded the priors provided by Boundary Guidance.
@@
Figure 6: Qualitative Analysis of Box-Focused Sampling for 3DGS Blobs. Box-Focused Sampling significantly retains more object blobs and reduces noisy background blobs. Note that we visualize only the positions of the Gaussian blobs to highlight their spatial distribution, while omitting other attributes for clarity.
[…]
Figure 7: Failure cases. In scenes with severe occlusions, our method exhibits a small number of missed detections. These failure cases are caused by occlusions, which introduce inherent ambiguity in multi-view detection when key visual evidence is obscured.